# PersonaMark: Personalized LLM watermarking for model protection and user attribution

### v1

@@ 5 Experiments
$\operatorname{PPL}(T)=P(t_{1},t_{2},t_{N})^{-\frac{1}{N}}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(t_{i}\mid t_{1},t_{2},t_{i-1})\right)$ (3)
It can be calculated by the Nth root of the reciprocal of the joint probability of the sentences in the test set, where N is the average number of words in the sentence. Where T represents a set of text sequences, including tokens such as t1, t2 to tN, and the probability in the perplexity formula can be calculated by a language model. The aspect based sentiment polarity is calculated using the deberta-v3-base-absa-v1.1 model333debertav3baseabsav1.1:https://huggingface.co/yangheng/deberta-v3-base-absa-v1.1. We utilize aspect based sentiment analysis rather than normal sentiment analysis since the prompts in the experiment are open-end generation task. Normal sentiment analysis tends to have greater fluctuations. We set the subjects of sentences as aspects and calculate the sentiment score according to the subjects. In this way, the average sentiment polarity of the model is determined as a metrics indicating the behavior of the models. The key information consistency is represented by the similarity obtained from the facebook/bart-large-cnn model444facebook/bart-large-cnn:https://huggingface.co/facebook/bart-large-cnnafter summarization using the Sentience Transformer (all-MiniLM-L6-v2) model [^5] Sentence Transformers documentation, https://www.sbert.net/docs/sentence_transformer/pretrained_models.html. For a given prompt, our approach’s aim is to let the models generate the same key components as the models without watermark, though the responded sentences might be different. In addition, the reading ease indicator is used to assist in evaluating the behavioral changes of models when generating text. The reading ease score is based on the Flesch Reading Ease Score (FRES) [^6] Flesch Reading Ease and Flesch Kincaid Grade Level:https://readable.com/readability/flesch-reading-ease-flesch-kincaid-grade-level/, which is used to evaluate the readability of an article or sentence from the perspective of the age required to read the sentence. For the English language, the formula for calculating sentence difficulty is as follows:
@@
We select the most commonly used watermarking method KGW as the baseline method. The KGW method is based on manipulating the probability of logits, dividing the vocabulary into green and red word lists with a certain degree of randomness. The logits corresponding to the red word list are adjusted to 0 or a penalty score is applied to make them appear less often. Finally, statistical tests are used to determine whether any given text has this specificity. Due to its concise design and model adaptability, it has become a commonly used baseline in the field of text watermarking, and lots of current research on text watermarking work on it for improvement. We measure the interference of watermarking methods on model generation behavior through a comprehensive evaluation of generated text, and combines text correctness indicator to control the degree of synonym attacks within a reasonable range. The text correctness is calculated by LanguageTool [^7] LanguageTool, style and grammar checker for 25+ languages, https://github.com/languagetool-org/languagetool. When the proportion of synonym attacks reaches 40%, through manual evaluation and GPT-4 based quality evaluation, it is found that the text already contains too many syntactic, lexical, and other errors. Therefore, the proportion of synonym is limited to 40% to fit real-world scenarios. We select four open-source model series as the base model for evaluation including Gemma-2B, Mistral-7B, OLMo-7B and Phi-3.5-mini. In order to conduct experimental comparisons under resource constraints, the random sampling strategy is used as an auxiliary supplementary method. When a smaller beam width causes the model to fall into a local minimum, disturbance intervention is performed by increasing randomness.
@@
Figure 3: AI-generated text detection performance on content generated by the PersonaMark
We also compare PersonaMark with several AI-generated text detectors to explore its ability under AI-generated text detection scenarios. Results show that our approach retains the ability for AI-generated detection task. In terms of the effectiveness of text detection generated by artificial intelligence, the F1 score of the detection results has significantly improved compared to the seven commonly used detection methods.
@@ 5.1 Further Discussions
Watermark Strength Analysis. As shown in Table 1, Figure 4, and Figure 5, KGW’s watermark detection performace on Phi-3.5 model is lower than other models. And when KGW is converted to KGW-Persona with a fixed green list to fit personalized settings, the AUC score drops much more. The reason behind this probably lies in the size of vocabulary. Because Phi-3.5 supports a vocabulary size of up to 32064 tokens, while other models’ vocabulary size is larger. For example, Gemma-2B-it has a vocabulary size of 256128. KGW and KGW-based methods inject their watermark by altering the probability of vocabulary. Restriced vocabulary size restrains the potential watermarking space, thus leading to worse performance and less robustness. This demonstrates our PersonaMark’s feasibility since sentences space is much larger than the fixed vocabulary of LLMs. The dropping performance might not comes from the number of params and size of models, since Gemma-2B is smaller than Phi-3.5 which is 3.8B.
@@
Comparing with KGW, the mainstream text watermarking baseline, PersonaMark achieved the best results in terms of perplexity, indicating that the text generated has the lowest interference on the text generation process. This comes from the level of watermarking, as KGW-based method manipulate the probability distribution of words. The probability of distribution is one of the goal of pretraining, and its modification introduces much bias. However, PersonaMark manipulates the distribution of sentences, whose candidate space is larger and the distribution inner sentences will not change. In terms of ABSA average sentiment polarity, semantic key information consistency, and sentence readability metrics, PersonaMark has the smallest difference compared to the text generated by models without watermarking. The baselines’ reading ease score drops more than PersonaMark. This may come from their watermarking scheme that part of vocabulary is randomly blocked as red list while common words are not evenly distributed in the vocabulary, which reduces the appearing probability of some common words and increases the appearing probability of some uncommon words. The increase of baselines in the frequency of uncommon words increases the difficulty of reading, while PersonaMark preserves more common words. In terms of key information consistency especially, the similarity values of PersonaMark is much better than the baseline method with the same base model and input instructions. This is not only comes from the level of watermarking, but also comes from the manipulation frequency. KGW will manipulate each word’s probability in a sentence while the manipulation frequency of PersonaMark is lower as each sentence is manipulated once. What’s more, the i-th word’s green list and red list division depends on previous words in the baselines, so the bias of introducing watermark is reinforced and accumulated sequentially, making the model further away from the original state.
@@
In terms of watermark recognition performance, the F1 and AUC values in this article are better than the baseline watermark model, with better classification performance and close to the theoretical maximum value. Rather than a non-interpretable deep learning detector, the Z-test of null and alternative hypotheses has statistical interpretability while improving the recognition performance of the detector. In terms of robustness, PersonaMark demonstrates high robustness against synonym attacks. This is due to the principle advantage of sentence structure representation strategy. When synonym attacks are limited by grammar and quality rules, they will not change the sentence structure representation, making it impossible to effectively attack our method.
@@
Table 2: Detection times of different user amounts.
During the watermark detection process, since the text might be generated with one of the hashing functions in the database. In order to determine exactly which hashing function matches the text or no one matches the text(means it is not watermarked by our approach), each user’s hashing function need to operate the hashing one time. For 100 users, the hashing process needs to repeat 100 times. When the number of users exceeds 1000, the detection speed increases linearly with the number of users. However, the detection speed is also fast due to the fast computation speed of hashing algorithms. As show in Table 2 it takes 3.86 seconds when there are 100,000 users. The detection speed could be faster with more CPU cores or with parallel computation.
@@ 5.2 Case Study
As shown in Table 6, the text generated by Gemma without watermark contains both bit-0 sentences and bit-1 sentences, denoted as grey and light blue respectively, the number of which should be roughly equal. Its Z-score is low and p-value is high. The null hypothesis will not be rejected and this piece of text will not be detected as watermarked. The text generated by Gemma applied PersonaMark algorithm contains only bit-1 sentences, denoted light blue. The Z-value is high and p-value is extremely low, which is enough to reject the null hypothesis. The text of PersonaMark is perturbed by large proportion of synonym attack to be Perturbed PersonaMark text. Only three bit-1 sentences is hashed as bit-0 after the synonym attacking perturbation, thus the Z-score and p-value almost remains and we can reject the null hypothesis with a high confidence(low p-value). The P value refers to the probability of a more extreme result than the sample observation when the null hypothesis is true. If the P value is very small, it means that the probability of the null hypothesis is very small. If it occurs, according to the principle of small probability, we have reason to reject the null hypothesis. The smaller the P value, the more reason we have to reject the null hypothesis.
@@
6 Conclusions
We employ sentence structure information as the medium for watermark injection. First, sentence structure representation is derived from syntactic analysis and we encode structural information using a carefully designed hash function. Next, utilizing a sentence-level beam search strategy, we sampled candidate sentences with encoding bit-1 information and successfully integrated the watermark into the generated text seamlessly. At the same time, by adding the user ID to the input of the hash function, we construct a hashing function database corresponding to specific users. Personalized hashing function provides personalized watermarking injection process. We achieve personalized watermarking by adding different watermarks for different users, which means the generated text under any user’s account will contain his unique user ID. Though watermark detection based on the hashing function database, the user ID could be retrieved.

### v2

@@ 5.1 Evaluation Metrics
$\begin{split}\operatorname{PPL}(T)&=P(t_{1},t_{2},\ldots,t_{N})^{-\frac{1}{N}}\\ &=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(t_{i}\mid t_{1},t_{2},\ldots,t_{i-1})\right)\end{split}$ (3)
It can be calculated by the Nth root of the reciprocal of the joint probability of the sentences in the test set, where N is the average number of words in the sentence. Where T represents a set of text sequences, including tokens such as t1, t2 to tN, and the probability in the perplexity formula can be calculated by a language model. The aspect based sentiment polarity is calculated using the deberta-v3-base-absa-v1.1 model333debertav3baseabsav1.1:https://huggingface.co/yangheng/deberta-v3-base-absa-v1.1. We utilize aspect based sentiment analysis rather than normal sentiment analysis since the prompts in the experiment are open-end generation task. Normal sentiment analysis tends to have greater fluctuations. We set the subjects of sentences as aspects and calculate the sentiment score according to the subjects. In this way, the average sentiment polarity of the model is determined as a metrics indicating the behavior of the models. The key information consistency is represented by the similarity obtained from the Sentience Transformer (all-MiniLM-L6-v2) model [^4] Sentence Transformers documentation:https://www.sbert.net/docs/sentence_transformer/pretrained_models.html after summarization using the facebook/bart-large-cnn model555facebook/bart-large-cnn:https://huggingface.co/facebook/bart-large-cnn. For a given prompt, our approach’s aim is to let the models generate the same key components as the models without watermark, though the responded sentences might be different. In addition, the reading ease indicator is used to assist in evaluating the behavioral changes of models when generating text. The reading ease score is based on the Flesch Reading Ease Score (FRES) [^6] Flesch Reading Ease and Flesch Kincaid Grade Level:https://readable.com/readability/flesch-reading-ease-flesch-kincaid-grade-level/, which is used to evaluate the readability of an article or sentence from the perspective of the age required to read the sentence. For the English language, the formula for calculating sentence difficulty is as follows:
@@
5.2 Baselines
We select the most commonly used watermarking method KGW as the baseline method. The KGW method is based on manipulating the probability of logits, dividing the vocabulary into green and red word lists with randomness. The logits corresponding to the red word list are adjusted to 0 or a penalty score is applied to make them appear less often. Finally, statistical tests are used to determine whether any given text has this specificity. Due to its concise design and model adaptability, it has become a commonly used baseline in the field of text watermarking, and lots of current research on text watermarking work on it for improvement. We measure the interference of watermarking methods on model generation behavior through a comprehensive evaluation of generated text, and combines text correctness indicator to control the degree of synonym attacks within a reasonable range. The text correctness is calculated by the following formula where C is the number of errors computed with LanguageTool [^7] LanguageTool, style and grammar checker for 25+ languages:https://github.com/languagetool-org/languagetool.
$\text{TextCorrectness}=\begin{cases}\log_{2}\left(\frac{1}{c}\right),&\text{if}C\neq 0\\ 1,&\text{if }C=0\end{cases}$ (5)
When the proportion of synonym attacks reaches 40%, through manual evaluation and GPT-4 based quality evaluation, it is found that the text already contains too many syntactic, lexical, and other errors. Therefore, the proportion of synonym is limited to 40% to fit real-world scenarios. We select four open-source model series as the base model for evaluation including Gemma-2B, Mistral-7B, OLMo-7B and Phi-3.5-mini. In order to conduct experimental comparisons under resource constraints, the random sampling strategy is used as an auxiliary supplementary method. When a smaller beam width causes the model to fall into a local minimum, disturbance intervention is performed by increasing randomness.
@@ 5.3 Platform and Generation Settings
Figure 3: AI-generated text detection performance on content generated by the PersonaMark. As watermarking techniques become one of the approaches for AI-generated text detection, we compare PersonaMark with five metric-based methods and a model-based method of this task[10]. Results show that our method aiming for multi-user attribution still retains the ability of zero-bit watermark on AI-generated text detection , with F1 value and AUC value close to the maximum value.
@@ 6.1 Effectiveness Analysis
Watermark Strength. As shown in Table 1, Figure 4, and Figure 5, KGW’s watermark detection performance on Phi-3.5 model is lower than other models. And when KGW is converted to KGW-Persona with a fixed green list to fit personalized settings, the AUC score drops significantly. This performance decline can likely be attributed to the vocabulary size: Phi-3.5 supports a vocabulary of up to 32,064 tokens, whereas other models, such as Gemma-2B-it, use a much larger vocabulary of 256,128 tokens. KGW and KGW-based methods inject their watermark by altering the token probability distribution. A restricted vocabulary size restrains the potential watermarking space, thus leading to worse performance and lower robustness. This finding supports the feasibility of PersonaMark, as its sentence-level watermarking operates in a far larger space than the fixed token vocabulary of LLMs. The observed performance drop is unlikely to be due to the number of parameters or model size, since Gemma-2B, which is smaller than Phi-3.5 (3.8B parameters), performs worse.
Model Behaviors. We measure model behavior with metrics describing features of the text outputs, including perplexity, ABSA sentiment polarity, reading ease score, and key information consistency. When compared to KGW, the mainstream text watermarking baseline, PersonaMark achieved the best results in terms of perplexity, indicating minimal interference with the text generation process. This comes from the level of watermarking, as KGW-based method manipulate the probability distribution of words. The probability of distribution is one of the goal of pretraining, and its modification introduces much bias. However, PersonaMark manipulates the sentence-level distribution, which is less intrusive because it operates over a larger space, and does not alter the internal distribution of tokens within sentences.
@@
In terms of ABSA average sentiment polarity, semantic key information consistency, and sentence readability, PersonaMark has the smallest difference compared to the text generated by models without watermarking. The reading ease score for baselines drops more than that for PersonaMark. This may come from their watermarking scheme that part of vocabulary is randomly blocked as red list while common words are not evenly distributed in the vocabulary, which reduces the appearing probability of some common words and increases the appearing probability of some uncommon words. The increase of baselines in the frequency of uncommon words increases the difficulty of reading, while PersonaMark preserves more common words. In terms of key information consistency especially, the similarity values of PersonaMark is much better than the baseline method with the same base model and input instructions. This is not only comes from the level of watermarking, but also comes from the manipulation frequency. KGW will manipulate each word’s probability in a sentence while the manipulation frequency of PersonaMark is lower as each sentence is manipulated once. What’s more, the i-th word’s green list and red list division depends on previous words in the baselines, so the bias of introducing watermark is reinforced and accumulated sequentially, making the model further away from the original state.
Table 2: Win rate of our PersonaMark on Fluency and Content Consistency
Models Fluency (%) Content Consistency (%)
Gemma-2B-IT 80 60
Mistral-7B-Instruct-V0.3 100 85
OLMo-7B-Instruct 75 65
Phi-3.5-mini-Instruct 100 100
Human Evaluation. We conduct a human evaluation by sampling 20 items from each model’s outputs, involving four evaluators comprising two doctoral students and two master students. The evaluation focuses on two dimensions: Fluency and content consistency. The evaluators are provided with the prompt used for generation, the original model output without watermark, the KGW-Persona output, and the PersonaMark output.. They are instructed to select the superior watermarked text based on its fluency and its consistency with the original output. The former dimension aims to assess the quality of texts, while the latter aims to indicate changes in model behavior when using identical prompts. The findings presented in Table 2 indicate that PersonaMark achieves significantly higher average win rates in both fluency and content consistency, with respective values of 88.75 and 77.75. These results markedly surpass those of the multi-user baseline, KGW-Persona. Notably, the Phi-3.5 model entirely fails to respond to the original prompt under KGW-Persona settings. For instance, it generates irrelevant content such as math, code, or biological knowledge for story generation prompts. In contrast, under our PersonaMark method, the model retains the ability to follow instructions promptly. The underlying cause for this discrepancy may be analogous to the rationale previously elucidated regarding watermark strength. The Phi-3.5 model exhibits a comparatively limited vocabulary size relative to other models, resulting in a residual vocabulary of only 16,032 tokens after deployment of the KGW-Persona watermarking. This substantial reduction in vocabulary severely impairs the model’s generative capabilities. This limitation underscores the inherent performance degradation of the KGW-like methods, attributable to the incongruity between logits manipulation and the training and inference processes of large language models.
@@
Watermark Recognition Performance. In terms of watermark recognition performance, the F1 and AUC values of our method are better than the baseline, with better classification performance and close to the theoretical maximum value. Rather than a non-interpretable deep learning detector, the Z-test of null and alternative hypotheses has statistical interpretability while improving the recognition performance of the detector. In terms of robustness, PersonaMark demonstrates high robustness against synonym attacks. This is due to the principle advantage of sentence structure representation strategy. When synonym attacks are limited by grammar and quality rules, they will not change the sentence structure representation, making it impossible to effectively attack our method.
Table 3: Detection time of different user amounts.
@@
Watermark Detection Speed. During the watermark detection process, since the text might be generated with one of the hashing functions in the database. In order to determine exactly which hashing function matches the text or no one matches the text(means it is not watermarked by our approach), each user’s hashing function need to operate the hashing one time. For 100 users, the hashing process needs to repeat 100 times. When the number of users exceeds 1000, the detection speed increases linearly with the number of users. However, the detection speed is also fast due to the fast computation speed of hashing algorithms. As show in Table 3 it takes 3.86 seconds when there are 100,000 users. The detection speed could be faster with more CPU cores or with parallel computation.
AI-generated text Detection. We also compare PersonaMark with several AI-generated text detectors to explore its ability under AI-generated text detection scenarios. Results show that our approach retains the ability for AI-generated detection task. In terms of the effectiveness of text detection generated by artificial intelligence, the F1 score of the detection results has significantly improved compared to the seven commonly used detection methods.
@@ 6.2 Case Study
As shown in Figure 6, the text generated by Gemma without watermark contains both bit-0 sentences and bit-1 sentences, denoted as grey and light blue respectively, the number of which should be roughly equal. Its Z-score is low and p-value is high. The null hypothesis will not be rejected and this piece of text will not be detected as watermarked. The text generated by Gemma applied PersonaMark algorithm contains only bit-1 sentences, denoted light blue. The Z-value is high and p-value is extremely low, which is enough to reject the null hypothesis. The text of PersonaMark is perturbed by large proportion of synonym attack to be Perturbed PersonaMark text. Only three bit-1 sentences is hashed as bit-0 after the synonym attacking perturbation, thus the Z-score and p-value almost remains and we can reject the null hypothesis with a high confidence(low p-value). The P value refers to the probability of a more extreme result than the sample observation when the null hypothesis is true. If the P value is very small, it means that the probability of the null hypothesis is very small. If it occurs, according to the principle of small probability, we have reason to reject the null hypothesis. The smaller the P value, the more reason we have to reject the null hypothesis.
@@
7 Discussions
7.1 Limitations
Since the watermark injection process leverages beam search, our method necessitates more computing resources and inference time compared to non-watermarking generation processes. While the method’s high scalability ensures that increased user numbers do not incur additional costs, the inference time for individual users is affected by the watermarking injection. Given that our method relies on sentence structures, further investigation might be required in contexts where sentence structures are less significant and do not convey stylistic features. Additionally, because PersonaMark operates at the sentence level, it is not applicable when users generate only a single sentence, representing a potential limitation. Nevertheless, in typical use cases, models are generally requested to produce outputs at the paragraph level or longer, which include multiple sentences. Moreover, even if one were to request single sentences repeatedly, the watermark would still be injected, thereby maintaining the method’s effectiveness.
7.2 Ethical Considerations
Our research on PersonaMark has been conducted with careful consideration of potential ethical implications across multiple dimensions.
Data Privacy. PersonaMark utilizes a personalized hashing function to generate unique watermarks for each user. To achieve this, we ensure that no personally identifiable information (PII) is collected, stored, or transmitted. The algorithms could be achieved with anonymous IDs without any user data.
Transparency. We have prioritized transparency by providing detailed documentation on how PersonaMark operates, including the methodologies for watermark embedding and detection. By making the watermarking process understandable, we empower users to make informed decisions about its implementation.
Minimizing Impact on Model Performance. A critical ethical consideration is ensuring that the introduction of watermarking does not degrade the performance of LLMs. Through extensive evaluations, we have demonstrated that PersonaMark maintains high text quality, preserves the natural behavior of the models. By optimizing the watermarking process to be subtle and minimally invasive, we mitigate potential negative impacts on user experience and model functionality.
Compliance with Legal and Ethical Standards. Our research complies with relevant data protection regulations, including the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA). We have conducted thorough reviews to ensure that PersonaMark adheres to legal standards concerning data privacy and user consent. Additionally, we follow ethical disclosure practices by transparently reporting our methodologies, findings, and any limitations associated with PersonaMark. This openness facilitates accountability and fosters trust within the research community and among end-users.
8 Conclusion
In this paper, we presented PersonaMark, a novel personalized text watermarking scheme designed to address critical issues of copyright protection, accountability, and traceability in the context of customized LLMs. We employ sentence structure information as the medium for watermark injection. First, sentence structure representation is derived from syntactic analysis and we encode structural information using a carefully designed hash function. Next, utilizing a sentence-level beam search strategy, we sampled candidate sentences with encoding bit-1 information and successfully integrated the watermark into the generated text seamlessly. At the same time, by adding the user ID to the input of the hash function, we construct a hashing function database corresponding to specific users. Personalized hashing function provides personalized watermarking injection process. We achieve personalized watermarking by adding different watermarks for different users, which means the generated text under any user’s account will contain his unique user ID. Though watermark detection based on the hashing function database, the user ID could be retrieved.